OpenStudy (anonymous):

Solve by completing the square. Your answers will involve a, b, or c. x2+bx+c=0

terenzreignz (terenzreignz):

There is no a in this question. But that's a technicality, I guess... You know how to complete the square?

hartnn (hartnn):

or is the question, \(ax^2+bx+c=0\) ??

Directrix (directrix):

x^2 + bx +c = 0 x^2 + bx = - c x^2 + bx + b^2/4 = -c + b^2/4 x^2 + bx + (b/2)^2 = -4c/4 + b^2/4 x^2 + bx + (b/2)^2= (b^2 - 4c) /4 (x + b/2) ^2 = (b^2 - 4c) /4 Take square roots of both sides. Solve for x. Don't forget the plus or minus.